# Rethinking Causal Discovery Through the Lens of Exchangeability

### v1

@@ B.2 Examples from the Synthetic Dataset
Figure 4: This Figure displays 32 examples randomly sampled from the developed synthetic dataset.
@@ Appendix C SynthNN specifications
In this appendix, we provide the full technical description of the convolutional neural network (CNN) used for binary classification, the image‐conversion and labeling procedures, and relevant implementation details omitted from the main text.
@@ C.1 Data preprocessing
The data is originally presented as two paired vectors ($x$ and $y$). This is converted into an image using the following steps:
@@
    Min–maxscale both $\mathbf{x}$ and $\mathbf{y}$ to $[0,1]$.
@@
    $i=\lfloor x\,(N-1)\rfloor,\quad j=\lfloor y\,(N-1)\rfloor.$
@@
    Initialize an $N\times N$ zero matrix and set $I[j,i]=1$ for each point $(i,j)$, where $N=50$.
Afterwards, since there is no inherent true direction, half the images must be assigned to the $Y\xrightarrow{}X$ direction, and their axis must be switched. This is required to ensure the network sees $50\%$ of $X\xrightarrow{}Y$ and $50\%$ of $Y\xrightarrow{}X$, so that it has no prior bias. Since the examples are drawn from different combinations of design choices (see Subsection 4.3), the examples within each set of design choices were guaranteed to have half of each class, to ensure the network does not learn any biases towards the design choices (and not the causal direction). Finally, in order to improve the results, a Gaussian filter is applied to the images with $\sigma=0.5$, exclusively during training.

### v2

@@ B.2 Examples from the Synthetic Dataset
Figure 5: 32 examples randomly sampled from the developed synthetic dataset.
@@ Appendix C SynthNN specifications
In this appendix, we provide the full technical description of the convolutional neural network (CNN) used for binary classification, including data preprocessing, labeling procedures, and relevant implementation details omitted from the main text.
Additionally, in bivariate causal discovery datasets, even though the cause and effect are well-defined, there is an inherent arbitrariness in the labeling process (which variable should be considered $X$, which $Y$, and consequently, the true label $X\to Y$ or $Y\to X$). Therefore, to ensure an unbiased training of the developed neural network, we must ensure a balanced weighted training dataset (with $50\%$ $X\xrightarrow{}Y$ and $50\%$ $Y\xrightarrow{}X$).
@@ C.1 Data preprocessing
The data is originally presented as two paired vectors ($X$ and $Y$). This is converted into a binary 50x50 image ($N=50$) using the following steps:
@@
    Min–max scale both $X$ and $Y$ to $[0,1]$.
@@
    $i_{n}=\lfloor X_{n}\,(N-1)\rfloor,\quad j_{n}=\lfloor Y_{n}\,(N-1)\rfloor.$
@@
    Initialize an $N\times N$ zero matrix and set $I[j,i]=1$ for each obtained point $(i_{n},j_{n})$.
Finally, during training, a Gaussian filter is applied to the images with $\sigma=0.5$.